OpenStudy (anonymous):

Mandana is ordering new tile for her kitchen floor. The area of the floor is 114.7 square feet, and she wants to order between 15% and 20% extra of the materials. Which of the following is a reasonable amount of tile for Mandana to order? A. 131 square feet B. 136 square feet C. 128 square feet D. 139 square feet

OpenStudy (anonymous):

Oprah makes $156.00 per day plus an additional 7% commission on her sales. Which equation could be used to find how much Oprah made in one day, r, if she sold x dollars worth? A. 0.07x + $156.00 = r B. 156x - $7.00 = r C. 7x + $156.00 = r D. 0.07x - $156.00 = r

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

o.o Does that say oprah makes 156 a day??? wow lucky...

OpenStudy (anonymous):

can u plz help me plz i neeed help u r the only one i know that will help

undeadknight26 (undeadknight26):

Oprah's question would be A i believe...still working on the first one...

undeadknight26 (undeadknight26):

First one i would this would be B but not sure...

OpenStudy (anonymous):

A survey was conducted on the salaries of 20 randomly selected college graduates with degrees in the same subject area. Each person surveyed graduated within the same 5-year period. Ten of the people surveyed attended a private university, while the other 10 people surveyed attended a public university of roughly the same size. Private Public First Quartile $58,000 $35,000 Second Quartile (Median) $71,000 $42,000 Third Quartile $85,000 $54,000 Based on the samples, what generalization can be made? A. The interquartile range for private universities is $27,000 more than for public universities. B. The top twenty-five percent of both the private and public university graduates surveyed earned more than $50,000 annually. C. The interquartile range for public universities is $19,000 more than for private universities. D. The top ten percent of both the private and public university graduates surveyed earned more than $60,000 annually.

OpenStudy (anonymous):

Amy and Stephanie were playing a number game. Amy picked a number and told Stephanie that if she added 7 to the number and multiplied the sum by 4, it would equal 8. Which equation could Stephanie use to find the number, x? A. 7(x + 4) = 8 B. 4(x + 7) = 8 C. 4(x) + 7 = 8 D. 4 + (x + 7) = 8

undeadknight26 (undeadknight26):

1 question at a time please...

OpenStudy (anonymous):

ok sorry

undeadknight26 (undeadknight26):

I think; A survey was conducted on the salaries of 20...is either C or A...not sure...so im not much help on these questions...Still working on the last one...

OpenStudy (anonymous):

is it b?

undeadknight26 (undeadknight26):

Yup...sorry i went afk...